# Asymptotic Inference in a Stationary Quantum Time Series

### v1

@@ 1.7 Main theorems
$\mathcal{E}_{n}\left(\Theta\right):=\left\{\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta\right\}$ (1.28)
where $A_{n}\left(a\right)$ is the $n\times n$ symbol matrix pertaining to $a$. Define also a corresponding classical geometric regression experiment $\mathcal{G}_{n}\left(\Theta\right)$ as follows. For any function $a\in\Theta$ define a set of functionals (local averages on $\left[-\pi,\pi\right]$) as
@@
$\mathcal{F}_{n}\left(\Theta\right):=\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{Geo}\left(p\left(J_{j,n}\left(a\right)\right)\right),a\in\Theta\right\}$ (1.30)
Consider the set of quantum spectral densities $a$ defined as follows: the set of real functions on $\left[-\pi,\pi\right]$, such that for some $\alpha>0$, $M>1$
@@
$\delta\left(\mathcal{E}_{n}\left(\Theta\right),\mathcal{F}_{n}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty,$
i.e. $\mathcal{F}_{n}\left(\Theta\right)$ is asymptotically more informative than $\mathcal{E}_{n}\left(\Theta\right)$: $\mathcal{E}_{n}\left(\Theta\right)\lesssim\mathcal{F}_{n}\left(\Theta\right)$.
@@
$dY_{\omega}=\mathrm{arc\cosh}\left(a\left(\omega\right)\right)d\omega+\left(2\pi/n\right)^{1/2}dW_{\omega}\text{, }\omega\in\left[-\pi,\pi\right].$ (1.33)
and where $dW_{\omega},$ $\omega\in\left[-\pi,\pi\right]$ is Gaussian white noise and $Y_{\omega}=\int_{-\pi}^{\omega}dY_{\omega}$. Here $Q_{n}\left(a\right)$ is a distribution on the measurable space $\left(C_{\left[-\pi,\pi\right]},\mathcal{B}\left(C_{\left[-\pi,\pi\right]}\right)\right)$. For $\Theta=\Theta_{1}\left(\alpha,M\right)$ consider the experiment $\mathcal{G}_{n}\left(\Theta\right)=\left\{Q_{n}\left(a\right),\;a\in\Theta\right\}$.
@@
i.e. $\mathcal{F}_{n}\left(\Theta\right)$ and $\mathcal{G}_{n}\left(\Theta\right)$ are asymptotically equivalent: $\mathcal{F}_{n}\left(\Theta\right)\approx\mathcal{G}_{n}\left(\Theta\right)$
This claim essentially follows from the results of [GN98]. It implies that for $\alpha>1$, for the quantum time series, the white noise model $\mathcal{G}_{n}^{\prime}\left(\Theta\right)$ is an upper information bound as well. Note that the function $\mathrm{arc\cosh}$ is the analog of the log-transformation of the spectral density in (1.5).
@@
i.e. $\mathcal{G}_{n}\left(\Theta\right)$ is asymptotically less informative than $\mathcal{E}_{n}\left(\Theta\right)$: $\mathcal{G}_{n}\left(\Theta\right)\lesssim\mathcal{E}_{n}\left(\Theta\right)$.
Clearly $\Theta_{2}\left(d,M\right)\subset\Theta_{1}\left(\alpha,M^{\prime}\right)$ for $\alpha=1$ and some $M^{\prime}>M$ (cf. Lemma 3.20 below). This implies
@@
Consider quantum statistical experiments $\mathcal{E}=\left\{\mathcal{A},\rho_{\theta},\theta\in\Theta\right\}$ and $\mathcal{F}:=\left\{\mathcal{B},\sigma_{\theta},\theta\in\Theta\right\}$ having the same parameter space. For the special case that $\mathcal{A}=\mathcal{B}$ define their trace norm distance
@@
$\mathcal{E}$ $\mathcal{\sim}$ $\mathcal{F}$ (equivalent): $\Delta\left(\mathcal{E},\mathcal{F}\right)=0$
$\mathcal{E}_{n}$ $\simeq$ $\mathcal{F}_{n}$ (asymptotically trace norm equivalent): $\Delta_{0}\left(\mathcal{F}_{n},\mathcal{E}_{n}\right)\rightarrow 0$
@@
$\mathcal{E}_{n}$ $\approx$ $\mathcal{F}_{n}$ (asymptotically equivalent): $\Delta\left(\mathcal{F}_{n},\mathcal{E}_{n}\right)\rightarrow 0$
Note that ”more informative” above is used in the sense of a semi-ordering, i.e. its actual meaning is ”at least as informative”. If $\mathcal{E},\mathcal{F}$ are classical experiments, where the trace norm distance is a multiple of the total variation distance between probability measures, the relation $\mathcal{E}_{n}\simeq\mathcal{F}_{n}\mathcal{\ }$will also be described as asymptotic equivalence in total variation.
@@ 2.2 Distance of states in terms of symbols
For general Gaussian states, explicit expression for $S\left(\rho||\sigma\right)$ in terms of the first two moments have been obtained ([PLOB17] and references therein). Below we give a special formula which focuses on the zero mean gauge invariant case, and writes out $S\left(\rho||\sigma\right)$ directly in terms of the symbols rather than the covariance matrices.
@@ 2.5 The geometric regression model
which corresponds to a one mode thermal state with covariance matrix $\frac{1}{2}\lambda I_{2}$, $\lambda>1$ [WPGP+12]. We obtain that $\mathcal{E}_{m}^{d}$ is equivalent to
@@ 3.5 A deficiency bound from limit distributions
We now show how uniform asymptotic normality an estimator can be used to establish a bound on the one sided Le Cam deficiency. The result is inspired by the two theorems in [M8̈0].
@@
Consider a sequence of experiments $\mathcal{P}_{n}=\left\{\Omega_{n},\mathcal{X}_{n},P_{n,\theta},\theta\in\Theta\right\}$ where $P_{n,\theta}$ are probability measures on $\left(\Omega_{n},\mathcal{X}_{n}\right)$ and $\Theta$ is a compact subset of $\mathbb{R}^{d}$. Assume that for a sequence of statistics $\hat{\theta}_{n}:\left(\Omega_{n},\mathcal{X}_{n}\right)\rightarrow\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ one has
@@
where the map $\theta\rightarrow\Sigma_{\theta}$ is continuous in the norm $\left\|\cdot\right\|_{2}$ for covariance matrices and $\Sigma_{\theta}>0,\theta\in\Theta$ . Then for experiments
@@
where $Z^{\prime}$ is a standard normal $d$-vector independent of $Y_{n}$. Now
$\left\|H_{\gamma}P_{n,\theta}^{\prime}-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}$ (3.96)
@@
Since the map $\theta\rightarrow\Sigma_{\theta}$ is continuous and $\Theta\subset\mathbb{R}^{d}$ is compact, the set $\left\{\Sigma_{\theta},\theta\in\Theta\right\}$ is compact in Hilbert-Schmidt norm. Then $\Sigma_{\theta}>0,\theta\in\Theta$ implies that (analogously to (A.55))
@@ 3.6 Le Cam’s globalization method
The ”heteroskedastic normal experiment” (3.92) resulting from Theorem 3.15 arises as a global approximation, roughly speaking, in regular parametric models with asymptotic normalized information matrix $\Sigma_{\theta}^{-1}$; cf. [LC75] and discussions in [Mam86], [Nus96]. We will utilize this result as a tool in our quest for lower information bounds for the quantum time series. Below we cite Le Cam’s original result and then give an application in our context.
@@
(Theorem 1 in [LC75]) Let $\mathcal{P}=\left\{P_{\theta},\theta\in\Theta\right\}$ and $\mathcal{Q}=\left\{Q_{\theta},\theta\in\Theta\right\}$ be two dominated experiments indexed by the set $\Theta$. Assume that $\Theta$ is metrized by $W$, that $0\leq a<b$ are given. Assume also that
@@ A.3 Uniform convergence in distribution
Let us define uniform convergence in distribution, following [IH81], Appendix I. Consider a sample space $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$; convergence in distribution of a sequence of probability measures $Q_{n}$ to some $Q$ is written $Q_{n}\Longrightarrow_{d}Q$. Assume on $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ there is a sequence of families of probability measures $\mathcal{P}_{n}=\left\{P_{n,\theta}\text{, }\theta\in\Theta\right\}$, $n\in\mathbb{N}$ where $\Theta$ is an arbitrary set. The family $\mathcal{P}_{n}$ is said to uniformly converge in distribution to a family $\mathcal{P}=\left\{P_{\theta}\text{, }\theta\in\Theta\right\}$ if for every bounded continuous function $g$ on $\mathbb{R}^{d}$ we have
@@
$\beta\left(P,Q\right):=\sup\left\{\left|\int f\left(dP-dQ\right)\right|:\left\|f\right\|_{BL}\leq 1\right\}.$
@@
$\left\|P-Q\right\|_{TV}=\sup_{A\in\mathfrak{B}_{\mathcal{X}}}\left|P\left(A\right)-Q\left(A\right)\right|.$
@@
Recall that the total variation metric satisfies
$\left\|P-Q\right\|_{TV}=\frac{1}{2}\sup\left\{\left|\int f\left(dP-dQ\right)\right|:\left\|f\right\|_{\infty}\leq 1\text{, }f\text{ measurable}\right\},$ (A.49)
hence
$\beta\left(P,Q\right)\leq 2\left\|P-Q\right\|_{TV}.$
@@
The latter relation means $P_{n,\theta_{n}}\rightarrow_{d}P_{\theta}$ along $n\in\mathcal{N}_{1}$, hence $\beta\left(P_{n,\theta_{n}},P_{\theta}\right)\rightarrow 0$, which contradicts (A.50).
@@
Define a sequence $\theta_{n}^{\ast}$, $n\in\mathbb{N}$ by $\theta_{n}^{\ast}=\theta_{n}$ for $n\in\mathcal{N}_{2}$, $\theta_{n}^{\ast}=\theta$ for $n\notin\mathcal{N}_{2}$. Then $\theta_{n}^{\ast}\rightarrow\theta$ and by (iii) we have $P_{\theta_{n}^{\ast}}\overset{d}{\implies}P_{\theta}$, which contradicts (A.51).
@@
and it is easy to see that (A.53) implies
$s_{1}=\inf_{\Sigma\in\mathcal{S}}\lambda_{\min}\left(\Sigma\right)>0.$ (A.55)
Then Lemma A.6 implies the claim.
@@ A.4 Geometric distribution
The fourth central moment of $X$ is, for r=1pr=1-p [Wei]

### v2

@@ 1.7 Main theorems
$\mathcal{E}_{n}\left(\Theta\right):=\left(\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta\right)$ (1.28)
where $A_{n}\left(a\right)$ is the $n\times n$ symbol matrix pertaining to $a$. Define also a corresponding classical geometric regression experiment $\mathcal{F}_{n}\left(\Theta\right)$ as follows. For any function $a\in\Theta$ define a set of functionals (local averages on $\left[-\pi,\pi\right]$) as
@@
$\mathcal{F}_{n}\left(\Theta\right):=\left({\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{Geo}\left(p\left(J_{j,n}\left(a\right)\right)\right),a\in\Theta\right).$ (1.30)
Consider the set $\Theta_{1}\left(\alpha,M\right)$ of quantum spectral densities $a$ defined as the set of real functions on $\left[-\pi,\pi\right]$, such that for some $\alpha>0$, $M>1$
@@
$\delta\left(\mathcal{F}_{n}\left(\Theta\right),\mathcal{E}_{n}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty,$
i.e. $\mathcal{F}_{n}\left(\Theta\right)$ is asymptotically more informative than $\mathcal{E}_{n}\left(\Theta\right)$: $\mathcal{E}_{n}\left(\Theta\right)\precsim\mathcal{F}_{n}\left(\Theta\right)$.
@@
$dY_{\omega}=\mathrm{arc\cosh}\left(a\left(\omega\right)\right)d\omega+\left(2\pi/n\right)^{1/2}dW_{\omega}\text{, }\omega\in\left[-\pi,\pi\right]$ (1.33)
where $dW_{\omega},$ $\omega\in\left[-\pi,\pi\right]$ is Gaussian white noise. Here $Q_{n}\left(a\right)$ is a distribution on the measurable space $\left(C_{\left[-\pi,\pi\right]},\mathcal{B}_{C_{\left[-\pi,\pi\right]}}\right)$ where $\mathcal{B}_{C_{\left[-\pi,\pi\right]}}$ is the pertaining Borel sigma-algebra. For $\Theta=\Theta_{1}\left(\alpha,M\right)$ consider the experiment $\mathcal{G}_{n}\left(\Theta\right)=\left(Q_{n}\left(a\right),\;a\in\Theta\right)$.
@@
i.e. $\mathcal{F}_{n}\left(\Theta\right)$ and $\mathcal{G}_{n}\left(\Theta\right)$ are asymptotically equivalent: $\mathcal{F}_{n}\left(\Theta\right)\approx\mathcal{G}_{n}\left(\Theta\right)$.
This claim essentially follows from the results of [GN98]. It implies that for $\alpha>1$, for the quantum time series, the white noise model $\mathcal{G}_{n}\left(\Theta\right)$ is an upper information bound as well. Note that the function $\mathrm{arc\cosh}$ is the analog of the log-transformation of the spectral density in (1.5).
@@
i.e. $\mathcal{E}_{n}\left(\Theta\right)$ is asymptotically more informative than $\mathcal{G}_{n}\left(\Theta\right)$: $\mathcal{G}_{n}\left(\Theta\right)\precsim\mathcal{E}_{n}\left(\Theta\right)$.
It is easy to see that for $\alpha>0$, one has $\Theta_{2}\left(d,M\right)\subset\Theta_{1}\left(\alpha,M^{\prime}\right)$ for $M^{\prime}=M\min\left(1,d^{2\alpha}\right)$. In view of Theorems 1.1 and 1.2 this implies
@@
Consider quantum statistical experiments $\mathcal{E}=\left\{\mathcal{A},\rho_{\theta},\theta\in\Theta\right\}$ and $\mathcal{F}:=\left\{\mathcal{B},\sigma_{\theta},\theta\in\Theta\right\}$ having the same parameter space. For the special case that $\mathcal{A}=\mathcal{B}$ define their predual norm distance
@@
$\mathcal{E}$ $\mathcal{\sim}$ $\mathcal{F}$ (equivalent): $\Delta\left(\mathcal{F},\mathcal{E}\right)=0$
$\mathcal{E}_{n}$ $\simeq$ $\mathcal{F}_{n}$ (asymptotically norm equivalent): $\Delta_{0}\left(\mathcal{F}_{n},\mathcal{E}_{n}\right)\rightarrow 0$
@@
$\mathcal{E}_{n}$ $\approx$ $\mathcal{F}_{n}$ (asymptotically equivalent): $\Delta\left(\mathcal{F}_{n},\mathcal{E}_{n}\right)\rightarrow 0.$
Note that ”more informative” above is used in the sense of a semi-ordering, i.e. its actual meaning is ”at least as informative”. If $\mathcal{E},\mathcal{F}$ are classical experiments, where the predual norm distance is a multiple of the total variation distance between probability measures, the relation $\mathcal{E}_{n}\simeq\mathcal{F}_{n}\mathcal{\ }$will also be described as asymptotic equivalence in total variation.
@@ 2.2 Distance of states in terms of symbols
For general Gaussian states, explicit expression for $S\left(\rho||\sigma\right)$ in terms of the first two moments have been obtained ([PLO+17] and references therein). Below we give a special formula which focuses on the zero mean gauge invariant case, and writes out $S\left(\rho||\sigma\right)$ directly in terms of the symbols rather than the covariance matrices.
@@ 2.5 The geometric regression model
which corresponds to a one mode thermal state with covariance matrix $\frac{1}{2}\lambda I_{2}$, $\lambda>1$ [WPG+12]. We obtain that $\mathcal{E}_{m}^{d}$ is equivalent to
@@ 3.5 A deficiency bound from limit distributions
We now show how uniform asymptotic normality an estimator can be used to establish a bound on the one sided Le Cam deficiency. The result is inspired by the two theorems in [MÜL80]. For the definition of uniform convergence in distribution, of the bounded Lipschitz norm $\left\|\cdot\right\|_{BL}$ for functions and the bounded Lipschitz metric $\beta$ for probability measures cf. Section A.3.
@@
Consider a sequence of experiments $\mathcal{P}_{n}=\left\{\Omega_{n},\mathcal{X}_{n},P_{n,\theta},\theta\in\Theta\right\}$ where $P_{n,\theta}$ are probability measures on $\left(\Omega_{n},\mathcal{X}_{n}\right)$, and $\Theta$ is a compact subset of $\mathbb{R}^{d}$. Assume that for a sequence of statistics $\hat{\theta}_{n}:\left(\Omega_{n},\mathcal{X}_{n}\right)\rightarrow\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ one has
@@
where the map $\theta\rightarrow\Sigma_{\theta}$ is continuous in the norm $\left\|\cdot\right\|_{2}$ for covariance matrices and $\Sigma_{\theta}>0,\theta\in\Theta$ . Assume each experiment $\mathcal{P}_{n}$ is dominated by a sigma-finite measure. Then for experiments
@@
where $Z^{\prime}$ is a standard normal $d$-vector independent of $Y_{n}$. Now for the total variation metric (cf. (A.48)) we have
$\left\|H_{\gamma}P_{n,\theta}^{\prime}-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}$
@@
Since the map $\theta\rightarrow\Sigma_{\theta}$ is continuous and $\Theta\subset\mathbb{R}^{d}$ is compact, the set $\left\{\Sigma_{\theta},\theta\in\Theta\right\}$ is compact in Hilbert-Schmidt norm. Then $\Sigma_{\theta}>0,\theta\in\Theta$ implies
@@
We note that the assumption that the experiments $\mathcal{P}_{n}$ be dominated is used only to fit the quantum version (1.24) of the deficiency $\delta\left(\mathcal{P}_{n},\mathcal{Q}_{n}\right)$ which covers classical experiment only if these are dominated (cf. Section A.1.11, last paragraph)
@@ 3.6 Le Cam’s globalization method
The ”heteroskedastic normal experiment” (3.92) resulting from Theorem 3.15 arises as a global approximation, roughly speaking, in regular parametric models with asymptotic normalized information matrix $\Sigma_{\theta}^{-1}$; cf. [LE 75] and discussions in [MAM86], [NUS96]. We will utilize this result as a tool in our quest for lower information bounds for the quantum time series. Below we cite Le Cam’s original result and then give an application in our context.
@@
(Theorem 1 in [LE 75]) Let $\mathcal{P}=\left\{P_{\theta},\theta\in\Theta\right\}$ and $\mathcal{Q}=\left\{Q_{\theta},\theta\in\Theta\right\}$ be two dominated experiments indexed by the set $\Theta$. Assume that $\Theta$ is metrized by $W$, that $0\leq a<b$ are given. Assume also that
@@ A.3 Uniform convergence in distribution
Let us define uniform convergence in distribution, following [IH81], Appendix I. Consider a sample space $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ where $\mathfrak{B}^{d}$ is the Borel sigma-algebra; convergence in distribution of a sequence of probability measures $Q_{n}$ to some $Q$ is written $Q_{n}\Longrightarrow_{d}Q$. Assume on $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ there is a sequence of families of probability measures $\mathcal{P}_{n}=\left\{P_{n,\theta}\text{, }\theta\in\Theta\right\}$, $n\in\mathbb{N}$ where $\Theta$ is an arbitrary set. The family $\mathcal{P}_{n}$ is said to uniformly converge in distribution to a family $\mathcal{P}=\left\{P_{\theta}\text{, }\theta\in\Theta\right\}$ if for every bounded continuous function $g$ on $\mathbb{R}^{d}$ we have
@@
$\beta\left(P,Q\right):=\sup\left\{\left|\int f\left(dP-dQ\right)\right|:\left\|f\right\|_{BL}\leq 1\right\}.$ (A.47)
@@
$\left\|P-Q\right\|_{TV}=\sup_{A\in\mathfrak{B}^{d}}\left|P\left(A\right)-Q\left(A\right)\right|.$ (A.48)
@@
Recall that the total variation metric satisfies (A.49); hence
$\beta\left(P,Q\right)\leq 2\left\|P-Q\right\|_{TV}$
@@
The latter relation means $P_{n,\theta_{n}}\Longrightarrow_{d}P_{\theta}$ along $n\in\mathcal{N}_{1}$, hence $\beta\left(P_{n,\theta_{n}},P_{\theta}\right)\rightarrow 0$, which contradicts (A.51).
@@
Define a sequence $\theta_{n}^{\ast}$, $n\in\mathbb{N}$ by $\theta_{n}^{\ast}=\theta_{n}$ for $n\in\mathcal{N}_{2}$, $\theta_{n}^{\ast}=\theta$ for $n\notin\mathcal{N}_{2}$. Then $\theta_{n}^{\ast}\rightarrow\theta$ and by (iii) we have $P_{n,\theta_{n}^{\ast}}\overset{d}{\implies}P_{\theta}$, which contradicts (A.52).
@@
so that (A.54) and Lemma A.6 imply the claim.
@@ A.4 Geometric distribution
The fourth central moment of $X$ is, for r=1pr=1-p [49]